It takes a barber 15 minutes to serve one customer.  
a. What is the capacity of the barber expressed in customers per hour?   customers per hour
       
b. Assuming the demand for the barber is 2 customers per hour, what is the flow rate?   customers per hour
       
c. Assuming the demand for the barber is 2 customers per hour, what is the utilization?   percent
       
d. Assuming the demand for the barber is 2 customers per hour, what is the cycle time?   minutes per customer
Answer #1

1.       Capacity of barber expressed in customers per hour = 60/15 =4 customers per hour

2.       Demand for barber =2 customer per hour

Flow rate = 2 customers per hour

3.       Demand for barber =2 customer per hour

Utilization = flow rate/capacity = 2/4 = 0.5 or 50%

4.       Cycle time = 0.5* 60 min.

= 30 minutes per customer